# Theory of Unsupervised Super-Resolution Data Assimilation with Conditional Variational Autoencoders: Estimating Background Covariances via Super-Resolution

### v4

@@ 3.3 Implementation of DA using CVAE
The loss function for VAEs, including CVAEs, is generally an ELBO [7, 8, 9, 11, 12]. This can be understood from the ELBO formula [Eq. (7)]. To evaluate $l_{\text{Gauss}}$, the analysis $\bm{a}$ and its covariance $V$ are first computed from given $\bm{x}$ and $\bm{y}$; the observation vector $\bm{\hat{y}}$ [$=\mathsf{H}(\bm{\hat{z}})$] is then reconstructed from $\bm{a}$ and $V$ via Eq. (5). The value of $l_{\text{Gauss}}$ decreases as $\bm{\hat{y}}$ becomes closer to the input $\bm{y}$. The architecture that reconstructs the input is a characteristic of autoencoders [10]. The second and third terms in Eq. (7) are the conditions for the estimated true state $\bm{\hat{z}}$. To incorporate these conditions, we use a CVAE, which is trained by minimizing the loss function $l_{\text{Gauss}}$. This training is unsupervised because $l_{\text{Gauss}}$ does not require the ground truth for $\bm{z}$. This unsupervised learning is achieved through the use of the ELBO because $l_{\text{Gauss}}$ is derived from the ELBO in Eq. (2c).
@@
The covariance matrices $R$ and $B$ in $l_{\text{Gauss}}$ are treated as prescribed parameters (i.e., hyperparameters). Since it is difficult to determine many hyperparameters, these matrices are usually assumed to be proportional to identity matrices (i.e., having zero off-diagonal elements) [7, 8, 9, 11, 12]. However, the off-diagonal elements of $B$ are essential for the non-local inference in DA (Section 3.2). We adopt super-resolution (SR) to realize non-local inference using CVAEs, leading to a new theory of SRDA.
4 SRDA using the ELBO
4.1 Integration of SR into DA
@@
We modify the probability distributions in Eqs. (3a)–(3c) for SRDA:
$q\left(\bm{z}\mid\bm{x},\bm{y}\right)$ $=$ $\frac{1}{\sqrt{(2\pi)^{n_{z}}\det V(\bm{x},\bm{y})}}$ (10a)
$\times$ $\exp\Bigg{\{}-\frac{1}{2}\left\lVert\bm{z}-\mathsf{A}_{\text{HR}}(\bm{x},\bm{y})\right\rVert^{2}_{V(\bm{x},\bm{y})^{-1}}\Bigg{\}},$
@@
The ELBO for SRDA is obtained by substituting Eqs. (10a)–(10c) into Eq. (2c), as in Section 3.1:
$l_{\text{SRDA}}$ $=$ $\frac{1}{2r}\left\lVert\bm{y}-\mathsf{H}(\bm{\hat{z}})\right\rVert^{2}+\frac{1}{2b}\left\lVert\mathsf{A}_{\text{HR}}(\bm{x},\bm{y})-\mathsf{F}(\bm{x})\right\rVert^{2}$ (13)
$\quad+\frac{1}{2}\sum_{i=1}^{n_{z}}\left\{\frac{v_{i}(\bm{x},\bm{y})}{b}-\ln\left[\frac{v_{i}(\bm{x},\bm{y})}{b}\right]\right\},$
@@
Eqs. (13) and (14a) correspond to Eqs. (7) and (5), respectively, where $V$ in Eq. (14a) is defined by Eq. (12). As in Eq. (7), the Monte Carlo method (i.e., the reparameterization trick [8]) is applied in Eq. (13) with the sample size being 1. In the right-hand side of Eq. (14a), the arguments of $\bm{x}$ and $\bm{y}$ are omitted, and the definition of $\bm{a}_{\text{HR}}$ [Eq. (14b)] is used. This $\bm{a}_{\text{HR}}$ is obtained by super-resolving $\bm{a}_{\text{LR}}$. The first term of $l_{\text{SRDA}}$ represents the difference between the observation vector $\bm{y}$ and the estimated true state $\mathsf{H}(\bm{\hat{z}})$ in the observation space. The second term is the difference between the super-resolved analysis $\bm{a}_{\text{HR}}$ [$=\mathsf{A}_{\text{HR}}(\bm{x},\bm{y})$] and the super-resolved background $\mathsf{F}(\bm{x})$. The third term is a convex function of $v_{i}$ that takes the minimum when $v_{i}(\bm{x},\bm{y})=b$. The ELBO (i.e., $l_{\text{SRDA}}$) describes the balance between the observations and the super-resolved background state in estimating the HR true state.
4.2 Case of linear SR
The effects of non-locality of SR are examined separately for linear and nonlinear SR. For linear SR, we show that SR operators act as background error covariance matrices.
Linear SR operators were frequently used in SR studies, especially in theoretical studies, before the emergence of deep learning [47]. When generation processes of LR images from HR images are linear (e.g., subsampling or blurring), the inverse processes, namely the SR processes, are also linear [47]. An SR operator is denoted by a matrix $F$ of size $n_{z}\times n_{x}$ ($n_{z}>n_{x}$) that maps $n_{x}$-dimensional LR vectors into $n_{z}$-dimensional HR vectors.
To simplify the discussion, we take the variances $v_{i}$ ($i=1,\ldots,n_{z}$) to be constant and small values, as in Section 3.2. The estimated true state $\bm{\hat{z}}$ in Eq. (14a) is then approximated as $\bm{\hat{z}}=F\bm{a}_{\text{LR}}$. Further, the third term in Eq. (13) is ignored because it becomes constant.
@@ 4.3 Case of nonlinear SR
Eqs. (17a) and (17b) are similar to backpropagation formulas using mean squared errors as a loss function [48]. The differences from the observations (i.e., errors), $H\mathsf{F}(\bm{a}^{(n)})-\bm{y}$, propagate backward through the SR neural network $\mathsf{F}$. This backpropagation updates the current solution $\bm{a}^{(n)}$ (Fig. 4). Note that all parameters of $\mathsf{F}$ remain fixed because $\mathsf{F}$ is already trained. Since trained SR neural networks generally exhibit non-locality [30, 31, 32, 33], the backpropagation updates the analysis vector $\bm{a}$ from observations over distant grid points (Fig. 4). We demonstrate that a trained SR model actually exhibits this non-locality in Section 6.2. Equation (17b) indicates that the differences from the super-resolved background state, $\mathsf{F}(\bm{a}^{(n)})-\mathsf{F}(\bm{x})$, also induce updates through backpropagation. The ELBO (16) determines the balance between the updates from observation and background values.
Figure 4: Schematic of the non-local update by backpropagation through an SR neural network (i.e., a non-linear SR model). A large error (red) is propagated backward through yellow nodes in the hidden layers and updates the LR input values.
Minimizing Eq. (16) through backpropagation may be computationally expensive, although automatic differentiation in deep learning frameworks facilitates the calculation of $\partial\mathsf{F}/\partial\bm{a}^{(n)}$ in Eq. (17b). As in Section 3.3, we propose to solve the minimization problem using CVAEs.
4.4 Loss function for SR models
@@
where ||$\cdot$|| ⋅ | denotes the absolute value, $i$ specifies vector components, and $P$ is the projection matrix into the observation space ($\mathsf{H}=P$). Typically, $P$ consists of interpolation and/or subsampling operations. The right-hand side shows that $l_{\text{SR}}$ is an L1 norm, which is frequently used in SR studies [27, 28, 29, 41]. Backpropagation is feasible because $l_{\text{SR}}$ is differentiable with respect to $\mathsf{F}(\bm{x})$. The SR model $\mathsf{F}$ is trained using $\bm{x}$ as input and $\bm{y}$ as target. When the training data size is sufficiently large, the trained SR model infers the median of $\bm{y}$ conditioned on $\bm{x}$ [49]. This result suggests that the SR model becomes unbiased if the median of the observation error distribution is zero.
4.5 Implementation of SRDA using CVAE
An efficient minimization of $l_{\text{SRDA}}$ in Eq. (13) is possible with CVAEs, as in Section 3.3, as long as the statistical properties of $\bm{x}$ and $\bm{y}$ do not change significantly.
The CVAE for SRDA consists of an encoder and a decoder (Fig. 5). The difference from the CVAE in Fig. 2 is that the encoder incorporates an SR model and performs SR and DA simultaneously. We use the SR model in Section 4.4, which is trained with the loss function in Eq. (18). The encoder integrates the LR background $\bm{x}$ and the HR observations $\bm{y}$, estimating the HR analysis $\bm{a}_{\text{HR}}$ through computing the LR analysis $\bm{a}_{\text{LR}}$. This computation of $\bm{a}_{\text{LR}}$ [$=A_{\text{LR}}(\bm{x},\bm{y})$] emulates the result of iterative backpropagation through $\mathsf{F}$ (i.e., the result of minimizing $l_{\text{SRDA-NL}}$ in Section 4.3). The encoder also estimates the uncertainty of $\bm{a}_{\text{HR}}$ [$=\mathsf{F}(\bm{a}_{\text{LR}})$] by calculating the variance $V$ in Eq. (12). The decoder acts as the observation operator, as in Fig. 2. The CVAE is trained in an unsupervised manner using $l_{\text{SRDA}}$ as the loss function.
@@ 4.5 Implementation of SRDA using CVAE
Rigorously speaking, a different SR model should be used within the encoder because the SR model in Section 4.4 is trained to super-resolve the LR background $\bm{x}$, which differs from the LR analysis $\bm{a}_{\text{LR}}$. This difference in input may introduce bias in $\bm{a}_{\text{LR}}$. In contrast, the output $\bm{a}_{\text{HR}}$ [$=\mathsf{F}(\bm{a}_{\text{LR}})$] is expected to be less biased because the SR model is unbiased if the observation error distribution has a zero median (Section 4.4). Thus, $\bm{a}_{\text{HR}}$ is used as initial conditions for LR fluid models by resizing $\bm{a}_{\text{HR}}$ to the LR (Section 4.6). One solution for making $\bm{a}_{\text{LR}}$ less biased is to conduct fine-tuning of the SR model $\mathsf{F}$ during the training of the CVAE. For simplicity, this study does not use such an advanced learning method, which remains a future research topic; instead, the parameters of $\mathsf{F}$ are fixed during the CVAE training.
4.6 Sequential SRDA using the CVAE
We introduce a sequential SRDA method applicable to CFD simulations. The background state $\bm{x}$ is hereafter referred to as the forecast state because it is a prior forecast given by an LR fluid model. The forecast and assimilation are executed alternately, as in other statistical DA methods [1, 2]. This assimilation is performed using a trained CVAE’s encoder.
@@ 4.6 Sequential SRDA using the CVAE
5 Methods
We evaluate the proposed SRDA method through numerical experiments similar to observing-system simulation experiments (OSSEs) [50], which are commonly used in testing DA methods for the atmosphere and ocean. The ground truth (or the nature run) is generated by an HR fluid model, from which synthetic observations are computed. The SRDA employs an LR version of the fluid model, and the resultant forecast and analysis are compared with the ground truth.

### v5

@@ III.3 Implementation of DA using CVAE
The loss function for VAEs, including CVAEs, is generally an ELBO.Kingma and Welling (2019, 2014); Rezende, Mohamed, and Wierstra (2014); Kingma et al. (2014); Sohn, Yan, and Lee (2015) This can be understood from the ELBO formula [Eq. (7)] and naturally leads to the use of CVAEs rather than other models (e.g., generative adversarial networksCreswell et al. (2018)) that do not have ELBOs as their loss functions. To evaluate $l_{\text{Gauss}}$ in Eq. (7), the analysis $\bm{a}$ and its covariance $V$ are first computed from given $\bm{x}$ and $\bm{y}$; the observation vector $\bm{\hat{y}}$ [$=\mathsf{H}(\bm{\hat{z}})$] is then reconstructed from $\bm{a}$ and $V$ via Eq. (5). The value of $l_{\text{Gauss}}$ decreases as $\bm{\hat{y}}$ becomes closer to the input $\bm{y}$. The architecture that reconstructs the input is a characteristic of autoencoders.Dong et al. (2018) The second and third terms in Eq. (7) are the conditions for the estimated true state $\bm{\hat{z}}$. To incorporate these conditions, we use a CVAE, which is trained by minimizing the loss function $l_{\text{Gauss}}$. This training is unsupervised because $l_{\text{Gauss}}$ does not require the ground truth for $\bm{z}$. This unsupervised learning is achieved through the use of the ELBO because $l_{\text{Gauss}}$ is derived from the ELBO in Eq. (2c).
@@
The covariance matrices $R$ and $B$ in $l_{\text{Gauss}}$ are treated as prescribed parameters (i.e., hyperparameters). Since it is difficult to determine many hyperparameters, these matrices are usually assumed to be proportional to identity matrices (i.e., having zero off-diagonal elements).Kingma and Welling (2019, 2014); Rezende, Mohamed, and Wierstra (2014); Kingma et al. (2014); Sohn, Yan, and Lee (2015) However, the off-diagonal elements of $B$ are essential for the non-local inference in DA (Section III.2). We adopt super-resolution (SR) to realize non-local inference using CVAEs, leading to a new theory of SRDA.
IV SRDA using the ELBO
IV.1 Integration of SR into DA
@@
We make two assumptions to derive the loss function for SRDA: (A) the ELBO in Eq. (2c) and (B’) the following Gaussian distributions. The first assumption is the same as in the theory for DA in Section III, while the second assumption is a modification from Eqs. (3a)–(3c):
$q\left(\bm{z}\mid\bm{x},\bm{y}\right)$ $=$ $\frac{1}{\sqrt{(2\pi)^{n_{z}}\det V(\bm{x},\bm{y})}}\;\exp\Bigg{\{}-\frac{1}{2}\left\lVert\bm{z}-\mathsf{A}_{\text{HR}}(\bm{x},\bm{y})\right\rVert^{2}_{V(\bm{x},\bm{y})^{-1}}\Bigg{\}},$ (10a)
@@
The ELBO for SRDA is obtained by substituting Eqs. (10a)–(10c) into Eq. (2c), as in Section III.1:
$l_{\text{SRDA}}$ $=$ $\frac{1}{2r}\left\lVert\bm{y}-\mathsf{H}(\bm{\hat{z}})\right\rVert^{2}+\frac{1}{2b}\left\lVert\mathsf{A}_{\text{HR}}(\bm{x},\bm{y})-\mathsf{F}(\bm{x})\right\rVert^{2}+\frac{1}{2}\sum_{i=1}^{n_{z}}\left\{\frac{v_{i}(\bm{x},\bm{y})}{b}-\ln\left[\frac{v_{i}(\bm{x},\bm{y})}{b}\right]\right\},$ (13)
@@
Eqs. (13) and (14a) correspond to Eqs. (7) and (5), respectively, where $V$ in Eq. (14a) is defined by Eq. (12). As in Eq. (7), the Monte Carlo method (i.e., the reparameterization trick)Kingma and Welling (2014) is applied in Eq. (13) with the sample size being 1. In the right-hand side of Eq. (14a), the arguments of $\bm{x}$ and $\bm{y}$ are omitted, and the definition of $\bm{a}_{\text{HR}}$ [Eq. (14b)] is used. This $\bm{a}_{\text{HR}}$ is obtained by super-resolving $\bm{a}_{\text{LR}}$. The first term of $l_{\text{SRDA}}$ represents the difference between the observation vector $\bm{y}$ and the estimated true state $\mathsf{H}(\bm{\hat{z}})$ in the observation space. The second term is the difference between the super-resolved analysis $\bm{a}_{\text{HR}}$ [$=\mathsf{A}_{\text{HR}}(\bm{x},\bm{y})$] and the super-resolved background $\mathsf{F}(\bm{x})$. The third term is a convex function of $v_{i}$ that takes the minimum when $v_{i}(\bm{x},\bm{y})=b$. The ELBO (i.e., $l_{\text{SRDA}}$) describes the balance between the observations and the super-resolved background state in estimating the HR true state.
IV.2 Case of linear SR
The loss function in Eq. (13) is applicable to both linear and nonlinear SR. This section and the next (Section IV.3) are for interpreting this loss function through examining the effects of non-locality of SR in the linear and nonlinear regimes, respectively. Here, we show that linear SR operators act as background error covariance matrices.
Linear SR operators were frequently used in SR studies, especially in theoretical studies, before the emergence of deep learning.Park, Park, and Kang (2003) When generation processes of LR images from HR images are linear (e.g., subsampling or blurring), the inverse processes, namely the SR processes, are also linear.Park, Park, and Kang (2003) An SR operator is denoted by a matrix $F$ of size $n_{z}\times n_{x}$ ($n_{z}>n_{x}$) that maps $n_{x}$-dimensional LR vectors into $n_{z}$-dimensional HR vectors.
To simplify the discussion, we take the variances $v_{i}$ ($i=1,\ldots,n_{z}$) to be constant and small values, as in Section III.2. The estimated true state $\bm{\hat{z}}$ in Eq. (14a) is then approximated as $\bm{\hat{z}}=F\bm{a}_{\text{LR}}$. Further, the third term in Eq. (13) is ignored because it becomes constant.
@@ IV.3 Case of nonlinear SR
Eqs. (17a) and (17b) are similar to backpropagation formulas that use mean squared errors as a loss function.Goodfellow, Bengio, and Courville (2016) The differences from the observations (i.e., errors), $H\mathsf{F}(\bm{a}^{(n)})-\bm{y}$, propagate backward through the SR neural network $\mathsf{F}$. This backpropagation updates the current solution $\bm{a}^{(n)}$ to make $H\mathsf{F}(\bm{a}^{(n)})$ closer to $\bm{y}$ (Fig. 4). Note that all parameters of $\mathsf{F}$ remain fixed because $\mathsf{F}$ is already trained. Since trained SR neural networks generally exhibit non-locality,Dong et al. (2014); Gu and Dong (2021); Chen et al. (2023); Zhou et al. (2023) the backpropagation updates the analysis vector $\bm{a}$ using observations from distant grid points (Fig. 4). We demonstrate that a trained SR model actually exhibits this non-locality in Section VI.2. Equation (17b) indicates that the differences from the super-resolved background state, $\mathsf{F}(\bm{a}^{(n)})-\mathsf{F}(\bm{x})$, also induce updates through backpropagation. The ELBO [Eq. (16)] determines the balance between the updates from observation and background values.
Figure 4: Schematic of the non-local update by backpropagation through an SR neural network (i.e., a non-linear SR model). A large error (red), $H\mathsf{F}(\bm{a}^{(n)})-\bm{y}$, propagates backward through yellow nodes in the hidden layers and updates the LR input $\bm{a}^{(n)}$ to make $H\mathsf{F}(\bm{a}^{(n)})$ closer to $\bm{y}$. See text for details.
Minimizing Eq. (16) through backpropagation may be computationally expensive, although automatic differentiation in deep learning frameworks facilitates the calculation of $\partial\mathsf{F}/\partial\bm{a}^{(n)}$ in Eq. (17b). As in Section III.3, we propose to solve the minimization problem using CVAEs.
IV.4 Loss function for SR models
@@
where ||$\cdot$|| ⋅ | denotes the absolute value, $i$ specifies vector components, and $P$ is the projection matrix into the observation space ($\mathsf{H}=P$). Typically, $P$ consists of interpolation and/or subsampling operations. The right-hand side shows that $l_{\text{SR}}$ is an L1 norm, which is frequently used in SR studies.Ha et al. (2019); Anwar, Khan, and Barnes (2020); Lepcha et al. (2023); Fukami, Fukagata, and Taira (2023) Backpropagation is feasible because $l_{\text{SR}}$ is differentiable with respect to $\mathsf{F}(\bm{x})$. The SR model $\mathsf{F}$ is trained using $\bm{x}$ as input and $\bm{y}$ as target. When the training data size is sufficiently large, the trained SR model infers the median of $\bm{y}$ conditioned on $\bm{x}$.Hastie, Tibshirani, and Friedman (2009) This suggests that the SR model becomes unbiased if the median of the observation error distribution is zero.
IV.5 Implementation of SRDA using CVAE
An efficient minimization of $l_{\text{SRDA}}$ in Eq. (13) is possible with CVAEs, as in Section III.3, as long as the statistical properties of $\bm{x}$ and $\bm{y}$ do not change significantly.
The CVAE for SRDA consists of an encoder and a decoder (Fig. 5). The difference from the CVAE in Fig. 2 is that the encoder incorporates an SR model and performs SR and DA simultaneously. We use the SR model in Section IV.4, which is trained with the loss function in Eq. (18). The encoder integrates the LR background $\bm{x}$ and the HR observations $\bm{y}$, estimating the HR analysis $\bm{a}_{\text{HR}}$ through computing the LR analysis $\bm{a}_{\text{LR}}$. This computation of $\bm{a}_{\text{LR}}$ [$=A_{\text{LR}}(\bm{x},\bm{y})$] emulates the result of iterative backpropagation through $\mathsf{F}$ (i.e., the result of minimizing $l_{\text{SRDA-NL}}$ in Section IV.3). The encoder also estimates the uncertainty of $\bm{a}_{\text{HR}}$ [$=\mathsf{F}(\bm{a}_{\text{LR}})$] by calculating the variance $V$ in Eq. (12). The decoder acts as the observation operator, as in Fig. 2. The CVAE is trained in an unsupervised manner using $l_{\text{SRDA}}$ in Eq. (13) as the loss function.
@@ IV.5 Implementation of SRDA using CVAE
Rigorously speaking, a different SR model should be used within the encoder because the SR model in Section IV.4 is trained to super-resolve the LR background $\bm{x}$, which differs from the LR analysis $\bm{a}_{\text{LR}}$. This difference in input may introduce bias in $\bm{a}_{\text{LR}}$. In contrast, the output $\bm{a}_{\text{HR}}$ [$=\mathsf{F}(\bm{a}_{\text{LR}})$] is expected to be less biased because the SR model is unbiased if the observation error distribution has a zero median (Section IV.4). Thus, $\bm{a}_{\text{HR}}$ is used as initial conditions for LR fluid models by resizing $\bm{a}_{\text{HR}}$ to the LR (Section IV.6). One solution for making $\bm{a}_{\text{LR}}$ less biased is to conduct fine-tuning of the SR model $\mathsf{F}$ during the training of the CVAE. For simplicity, this study does not use such an advanced learning method, which remains a future research topic; instead, the parameters of $\mathsf{F}$ are fixed during the CVAE training.
IV.6 Sequential SRDA using the CVAE
We introduce a sequential SRDA method applicable to CFD simulations. The background state $\bm{x}$ is hereafter referred to as the forecast state because it is a prior forecast given by an LR fluid model. The forecast and assimilation are executed alternately, as in other statistical DA methods.Asch, Bocquet, and Nodet (2016); Carrassi et al. (2018) This assimilation is performed using a trained CVAE’s encoder.
@@ IV.6 Sequential SRDA using the CVAE
IV.7 Related studies
We first compare the present SRDA with previous SRDA methods.Barthélémy et al. (2022); Yasuda and Onishi (2023) The major differences are that the present SRDA is based on unsupervised learning and makes full use of the non-locality of SR. Barthélémy et al.Barthélémy et al. (2022) first integrated SR and DA, in which forecasts are super-resolved by a neural network trained via supervised learning, and then assimilated by an ensemble Kalman filter. Although this approach can be easily incorporated into existing DA frameworks due to the separation of SR and DA, the non-local nature of SR may not be fully exploited in the assimilation process. Our previous studyYasuda and Onishi (2023) proposed a different SRDA that performs end-to-end learning to jointly optimize SR and DA. This SRDA requires HR ground-truth data and does not provide a clear theoretical background for combining SR and DA. By contrast, the present SRDA does not rely on HR ground-truth data because it employs unsupervised learning based on the ELBO, and clarifies that the non-locality of SR benefits DA.
Other DA studies also use ELBOs.Ghimire et al. (2017); Xie et al. (2023); Lafon, Fablet, and Naveau (2023) The most distinct point from these studies is that they do not utilize VAEs. Indeed, the use of ELBOs is generally distinguished from the use of VAEs. When VAEs are used, ELBOs or underlying probability distributions often need to be designed specifically for VAEs.Kingma et al. (2014); Sohn, Yan, and Lee (2015) In our case, the approximate posterior $q\left(\bm{z}\mid\bm{x},\bm{y}\right)$ is conditioned on both $\bm{x}$ and $\bm{y}$ in Eq. (2c), which is essential for the encoder to perform assimilation (Sections III.3 and IV.5). Thus, one of our contributions is to provide a basis for the combination of ELBOs and VAEs for DA.
VAEs are already used for DA (see also Section I). Melinc and ZaplotnikMelinc, Boštjan and Zaplotnik, Žiga (2024) performed dimensionality reduction using a VAE and then applied the 3D-Var in the latent space. Glyn-Davies et al.Glyn-Davies et al. (2024) employed a dynamical VAE (DVAEGirin et al. (2021)) to encode time series of observational data, which were then assimilated using a Kalman filter. Xiao et al.Xiao et al. (2025) incorporated non-Gaussian background errors into the 3D-Var objective function by using the transformation with a VAE’s decoder. The key distinction of our approach is that the CVAE itself performs assimilation (i.e., integrating observations into forecasts) using the non-locality of SR. This integration can be understood as the update by backward propagation of discrepancies from observations (Section IV.3).
Fundamentally, both DA and deep learning can be formulated using backpropagation with adjoint equations.Abarbanel, Rozdeba, and Shirman (2018); Chen et al. (2018); Geer (2021) Using this analogy, recent studies have developed neural networks that update forecasts by propagating errors from observations along time.Lafon, Fablet, and Naveau (2023); Fablet et al. (2021); Beauchamp et al. (2022) Our results reveal that SR models can act as error propagators for DA along the spatial dimensions, implying the possibility that neural networks act as propagators over the four dimensions (i.e., three spatial and one time dimensions).
V Methods
We evaluate the proposed SRDA method through numerical experiments similar to observing-system simulation experiments (OSSEs),Fletcher (2023) which are commonly used in testing DA methods for the atmosphere and ocean. The ground truth (or the nature run) is generated by an HR fluid model, from which synthetic observations are computed. The SRDA employs an LR version of the fluid model, and the resultant forecast and analysis are compared with the ground truth.